OpenStudy (anonymous):

Perform the indicated operation. (-10) -(-4) _______ -14 -6 6 14

OpenStudy (shamil98):

-10 - (-4) -10 + 4 = -6 - x - = +

OpenStudy (anonymous):

@shamil98 thank you